OpenStudy (anonymous):

When you subtract 22 from a number and the result is doubled, the answer is 6 more than the original number. What is the number?

OpenStudy (millsemily):

2*(n - 22) = n + 6 2n - 44 = n + 6 n - 44 = 6 n = 50

OpenStudy (aripotta):

ok so we're going to say "a number" is x. "when you subtract 22 from a number" we can write as x - 22 "and the result is doubled" 2(x - 22) "the answer is 6 more than the original number" 2(x - 22) = x + 6

OpenStudy (aripotta):

now you can use 2(x - 22) = x + 6, and solve for x